PLEASE GIVE ME AN EXAMPLE!

  • HAVE to know: Dihybrid Crosses (2 genes), **KNOW the dihybrid F2 phenotypic ratio is 9:3:3:1 (9/16dom/dom, 3/16 dom/rec, 3/16 rec/dom, 1/16 rec/rec)

1.- HAVE to know how to make a dihybrid Punnetts quare
2.- HAVE to know how to do 2 separate monohybrid Punnett squares and use the Product Law & Fork-Line Method to find phenotypic probabilities.

Fork-Line Method to find phenotypic probabilities
Answer #1

Tall = TT, Tt

Short = tt

Purple flowers = PP, Pp

White flowers = pp

TTPP × ttpp = F1 = TtPp (tall and purple)

TtPp × TtPp = F2

Gametes TP Tp tP tp
TP TTPP TTPp TtPP TtPp
Tp TTPp TTpp TtPp Ttpp
tP TtPP TtPp ttPP ttPp
tp TtPp Ttpp ttPp ttpp

Phenotypic ratio = 9 : 3 : 3 : 1

9 = T_P_ = tall purple

3 = T_pp = tall white

3 = ttP_ = short purple

1 = ttpp = short white

By monohybrid crosses -

TtPp × TtPp

Tt × Tt

Gametes T t
T TT Tt
t Tt tt

PR = 3/4 (tall) : 1/4 (short)

Pp × Pp

Gametes P p
P PP Pp
p Pp pp

PR = 3/4 (purple) : 1/4 (white)

Probability of -

Tall purple = 3/4 × 3/4 = 9/16

Tall white = 3/4 × 1/4 = 3/16

Short purple = 1/4 × 3/4 = 3/16

Short white = 1/4 × 1/4 = 1/16

Fork line method -

Ttep x т+ Фр Tt X Tt Pp x Pp — А Р Т от теР – у | - 2 Рp — ттрр - yь | Rata Рp Tt PP = 2/16 3-7-tv. —> тер = 4/14 3 = te- Тер

Please rate high.